OpenStudy (anonymous):

Simplify the following expression. 7 - 4(5 + 3)2 a 1,536 b 135 c -249 d -4

OpenStudy (anonymous):

is that times 2 or squared ?

OpenStudy (anonymous):

squared

OpenStudy (anonymous):

Okay so 7-4=3

OpenStudy (anonymous):

5+3=8

OpenStudy (anonymous):

so now we have 3(8)^2

OpenStudy (anonymous):

do you know what's next ?

OpenStudy (anonymous):

3-8 (2)??

OpenStudy (anonymous):

where did you get 3-8 from we are multiplying here

OpenStudy (anonymous):

ohhh 3 x 8 = 24

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

24 squared is 576 but i don't see that as one of your choices

OpenStudy (anonymous):

and 64*3=192

OpenStudy (anonymous):

which isn't a choice either hmmm lemme ask @ganeshie8

OpenStudy (anonymous):

this is hard

OpenStudy (anonymous):

yea .. @tgawade

OpenStudy (anonymous):

First simplify ( 5+3)^2 we add both terms so (8)^2 = 8 * 8 = 64 okay ?

OpenStudy (anonymous):

So we have 7 - 4 (64)

OpenStudy (anonymous):

now multiply -4 * 64 = -256 So we have 7 - 256 = -249

OpenStudy (anonymous):

Do u get it ?